# Exploring Holomorphic Retracts

### v6

@@ Conclusion
We recall for one last time, that the class of homogeneous bounded balanced domains is precisely the class of bounded balanced symmetric domains and that any domain of this class, is biholomorphic to a Cartesian product of Cartan domains of types I – IV (the second chapter, specifically remark 2.3.8 of the book [Jrncki_invrnt_dst] and references therein). It follows that a domain $\Omega$ as in the above theorem is a polyball (and convex), in-particular.
5. Proof of Theorem 1.9
Before discussing our result as in the title of this section on the retracts of $\ell_{q}$ balls for $10<q<$ 1 we first review the case $q\geq 1$. The main theorem about this case was already mentioned in theorem 1.7. However that formulation is highly coordinate dependent. Here we first re-formulate it in-terms of commutavity with standard projections; the proof is based on an idea from the work [Baronti-Pappini] of Baronti – Pappini.
@@
$R$ is a $k$-dimensional retract of the $\ell^{p}$-ball in $\mathbb{C}^{N}$ if and only if one of the following holds:
  1. (1)
    There exists a norm-one projection which maps $\mathbb{B}_{p}$ onto $R$ which commutes with the standard orthogonal projection onto one of the $(N-k)$ or $(N-(k-1))$-dimensional subspace spanned by co-ordinate vectors.
  2. (2)
    There exist a finite sequence of projections $P_{k},P_{k-1},\ldots,P_{1}$ such that for each $r=1,2,\ldots,k$, $P_{r}$ is a projection of norm one which maps $\mathbb{C}^{N}$ onto an $r$-dimensional subspace and $P_{k}(\mathbb{B}_{p})=R$.
@@
It is easy to check that one implication of the theorem holds by using the definition of norm one projection. So we have only to prove the other implication. Assume that $R$ is a $k$-dimensional linear retract of $\mathbb{B}_{p}$. By theorem 2.a.4 in [Lindstraus], we obtain that $R=L\cap\mathbb{B}_{p}$ where $L$ is a $k$-dimensional linear subspace of $\mathbb{C}^{N}$ spanned by mutually disjointly supported vectors. Hence $L$ can be written as the solution set of $(N-k)$ linear equations, where each linear equation depends on atmost two variables. After applying Gaussian elimination and renaming variables, there is no loss of generality assume that above mentioned system of equations can be written as $AZ=0$, where $A$ is a $(N-k)\times N$ matrix $[B~{}B^{\prime}~{}B^{\prime\prime}]^{T}$ and $Z=[z_{1}~{}\ldots~{}z_{N}]^{T}$. The submatrices $B,B^{\prime},B^{\prime\prime}$ are of the following form:
  1. (i)
    $B=[B_{0}~{}B_{1}~{}\ldots B_{r-1}]^{T}$, where for each $i=0,1,\ldots,r-1$, $B_{i}=[b_{k_{i}+1}\ldots b_{k_{i+1}}]^{T}$ and $0=k_{0}<k_{1}<\ldots<k_{r}$. Futhermore for each $i$, $B_{i}$ satisfies the following properties:
    1. (a)
      For $j=k_{i}+1$ to $k_{i+1}-1$, $supp(b_{j})\cap supp(b_{j+1})\neq\phi$ and $|supp(b_{j})\cap supp(b_{j+1})|=1$.
    2. (b)
      For $i=0,1,\ldots,r-1$, if we define $supp(B_{i}):=\bigcup_{j=k_{i}+1}^{k_{i+1}}supp(b_{j})$ then $supp(B_{i})\cap supp(B_{j})=\phi$ for $i\neq j$.
  2. (ii)
    $B^{\prime}$ is of the form $[b_{1}^{\prime}~{}\ldots~{}b_{l}^{\prime}]^{T}$, where $supp(b_{i}^{\prime})\cap supp(b_{j}^{\prime})=\phi$ for $i,j=1,2,\ldots,l$ with $i\neq j$.
  3. (iii)
    $B^{\prime\prime}$ is of the form $[e_{s_{0}+1}~{}\ldots e_{s_{1}}]^{T}$. where $s_{0}=k_{r}+r+2l,s_{1}=N-k+r+l$ and $e_{i}$ denotes the standard basis of $\mathbb{C}^{N}$.
Note that $|supp(B)|=|\bigcup_{i=0}^{r-1}supp(B_{i})|=k_{r}+r$, $|supp(B^{\prime})|=2l$ and $|supp(B^{\prime\prime})|=s_{1}-s_{0}$. Since the number of equation is $(N-k)$, we have $k_{r}+l+(s_{1}-s_{0})=N-k$. Because of this equation, we take $s_{1}=N-k+r+l$. Hence, we obtain the total number of dependent variables in the system of equations $AZ=0$ is equal to
$|supp(B)|+|supp(B^{\prime})|+|supp(B^{\prime\prime})|$ $=$ $k_{r}+r+2l+(s_{1}-s_{0})$
$=$ $k_{r}+r+2l+(N-k)-(k_{r}+l)$
$=$ $N-k+r+l.$
First we want to construct mutually disjoint supported vectors $(v_{i})_{i=1}^{k}$ such that $(v_{i})$ forms a basis for $L$. For each $i=1,\ldots,r$, consider the system of equations $B_{i}Z_{i}=0$, where $Z_{i}=[z_{k_{i-1}+1}~{}\ldots~{}z_{k_{i}}]^{T}$. This solution space is spanned by a non-zero vector $w_{i}=(\mu_{k_{i-1}+1},\ldots\mu_{k_{i}})\in\mathbb{C}^{k_{i}-k_{i-1}}$. Since every two co-ordinates of $w_{i}$ satisfies a linear equation in two variables, every component of $w_{i}$ is non-zero. For each $i=1,\ldots,r$, define $v_{i}:=\mu_{k_{i-1}+1}e_{k_{i-1}+1}+\ldots+\mu_{k_{i}}e_{k_{i}}$, where $e_{i}$ denote the standard basis in $\mathbb{C}^{N}$. Note that $v_{i}\in L$ and $supp(v_{i})=k_{i}-k_{i-1}$. For $i=r+1$ to $r+l$, choose a non-zero vector $v_{i}$ which is orthogonal to $b^{\prime}_{i-r}$ and $supp(v_{i})=supp(b^{\prime}_{i-r})$. For $i=r+l+1$ to $k$, $v_{i}=e_{i+t}$, where $t=N-k$.
Define $u_{i}:=\frac{v_{i}}{\|v_{i}\|_{p}}$ for $i=1,2,\ldots,k$. Note that $(u_{i})_{i=1}^{k}$ forms a mutually disjointly supported vectors and a basis for $L$. Observe that $|supp(u_{i})|=2$ for $i=r+1$ to $r+l$ and $|supp(u_{i})|=1$ for $i=r+l+1$ to $k$. Using the idea in the proof of theorem 2.a.4 in [Lindstraus], we can construct projection $P$ of norm $1$ from given vectors $u_{1},\ldots,u_{k}$. Let $\sigma_{j}:=supp(u_{j})$ for $j=1,2,\ldots,k$. If we write $u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i}$, then we define the linear functional $u_{j}^{*}=\sum_{i\in\sigma_{j}}\frac{|\lambda_{i}|^{p}}{\lambda_{i}}e_{i}^{*}$. Observe that for $i=1,2,\ldots,r$, the linear functional $u_{i}^{*}$ depends only on the variables $z_{k_{i-1}+1},\ldots,z_{k_{i}}$, for $i=r+1,\ldots,r+l$, the linear functional $u_{i}^{*}$ depends only on the variables $z_{k_{r}+2i-r-1}$, $z_{k_{r}+2i-r}$ and for each $i\neq j$, $u_{j}^{*}$ does not depends on the variables $z_{k_{r}+2i-r-1}$ and $z_{k_{r}+2i-r}$. For $i=r+l+1$ to $k$, $u_{i}^{*}(z)=z_{i+t}$.
Define a projection $P:\mathbb{C}^{N}\to\mathbb{C}^{N}$ by
$P(z)=\sum_{j=1}^{k}u_{j}^{*}(z)u_{j}.$
Note that $P$ is a projection of norm $1$ onto $L$. Next we are looking for standard projections which commutes with $P$. For $i=1,\ldots,r$, consider the standard projection,
$\pi_{k_{i-1}+1,\ldots,k_{i}}(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$
Note that
$P\circ\pi_{k_{i-1}+1,\ldots,k_{i}}(z_{1},\ldots,z_{N})=P(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$
For each $i=1,\ldots,r$, note that $k_{i-1}+1,\ldots,k_{i}$ components of $P$ only depends on $z_{k_{i-1}+1},\ldots,z_{k_{i}}$. So, $\pi_{k_{i-1}+1,\ldots,k_{i}}\circ P(z_{1},\ldots,z_{N})=P(z_{1},\ldots,z_{k_{i-1}},0,\ldots,0,z_{k_{i}+1},\ldots,z_{N})$.
Let $m_{i}:=k_{r}+2i-r-1$. For $i=r+1,\ldots,r+l$, consider another projection
$\pi_{m_{i},m_{i}+1}(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{m_{i}-1},0,0,z_{m_{i+2}},\ldots,z_{N})$
Using similar arguments as mentioned above, we can conclude that $P\circ\pi_{m_{i},m_{i}+1}=\pi_{s_{i},s_{i}+1}\circ P$. As we noted before, the total number of dependent variables in the system of equation $AZ=0$ is $N-k+r+l$. So, we consider another projection $\pi$ defined by $\pi(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{s_{1}},0,\ldots,0)$. Note that $P$ commutes with $\pi$ and $rank(\pi)=N-k+r+l$. If $r+l\geq 1$, then rank of $\pi$ is atleast $N-(k-1)$. If $r+l=0$, then $rank(\pi)$ is atleast $(N-k)$. This finishes the proof of (i).
To prove (ii), take $P_{k}:=P$, which is a projection of norm one defined in the proof of (i). From the proof of (i), we obtain a standard projection $\pi_{k-1}$ of rank $(N-1)$ which commute with $P$. Let $P_{k-1}:=\pi_{k-1}\circ P$. Observe that $P_{k-1}$ is a norm one projection and $\rm image(P_{k-1})$ is a $k$-dimensional is a $(k-1)$-dimensional subspace of $\mathbb{C}^{N}$. Again applying the same argument to $P_{k-1}$, we get another projection $\pi_{k-2}$ of rank $(N-1)$ which commutes with $P$. Define $P_{k-2}:=\pi_{k-2}\circ P_{k-1}=\pi_{k-2}\circ\pi_{k-1}\circ P$. Hence we get that $P_{k-2}$ is a projection of norm one and it’s image is a $(k-2)$ dimensional subspace. Continuing this process, until we will obtain a projection $P_{1}$ of norm one such that $\rm image(P_{1})$ is a one dimensional subspace of $\mathbb{C}^{N}$. ∎
@@
An apparently even more geometric formulation of the retracts of $\ell_{p}$ balls for $1\leq p<\infty,p\neq 2$ is that the retracts are precisely isometric copies of lower dimensional $\ell_{p}$-balls; see theorem 6.3 of [Beata-survey]. This is a feature shared by the $\ell_{p}$-‘balls’ even for positive values of the parameter $p<1$. Indeed, we now move onto this case dealt with by theorem 1.9. Before going into its proof 1.9, we need to make a couple of observations about the geometry of the $\ell_{q}$ ball $D_{q}$ in $\mathbb{C}^{N}$ (for $q<1$). Firstly, although its absolute image of is concave (easily and vividly seen in the two-dimensional case), $D_{q}$ is actually not concave (nor convex) domain in $\mathbb{C}^{N}$; not even locally, near any of the smooth boundary points. We may establish this by computing the real Hessian at any smooth boundary point with positive real coordinates as $D_{q}$ is multicircular; we show this in the two-dimensional setting for simplicity, while it holds in higher dimensions as well by straight-forward extension of the arguments and computations. Rewrite the defining function for $D_{q}$ as
@@
$H$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial z\partial\overline{w}}\\ \frac{\partial^{2}\varphi}{\partial w\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial w\partial\overline{w}}\end{bmatrix}=\begin{bmatrix}\frac{q^{2}}{4}|z|^{q-2}&0\\ 0&\frac{q^{2}}{4}|w|^{q-2},\end{bmatrix}$
$\tilde{H}$ $=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z^{2}}&\frac{\partial^{2}\varphi}{\partial z\partial w}\\ \frac{\partial^{2}\varphi}{\partial w\partial z}&\frac{\partial^{2}\varphi}{\partial w^{2}}\end{bmatrix}=\begin{bmatrix}\frac{q}{2}(\frac{q}{2}-1)|z|^{q}z^{-2}&0\\ 0&\frac{q}{2}(\frac{q}{2}-1)|w|^{q}w^{-2}\end{bmatrix}$
@@
(5.1) $vH\overline{v}^{T}|_{(x,y)}+Re(vHv^{T})_{(x,y)}\\ =\frac{q^{2}}{4}\left(x^{q-2}|v_{1}|^{2}+y^{q-2}|v_{2}|^{2}\right)+\frac{q}{2}(\frac{q}{2}-1)Re\left(x^{q-2}v_{1}^{2}+y^{q-2}v_{2}^{2}\right).$

### v7

@@ Conclusion
We recall for one last time, that the class of homogeneous bounded it balanced domains is precisely the class of bounded balanced symmetric domains and that any domain of this class, is biholomorphic to a Cartesian product of Cartan domains of types I – IV (the second chapter, specifically remark 2.3.8 of the book [Jrncki_invrnt_dst] and references therein), a polyball.
5. Proof of Theorems 1.8 and 1.10
Before discussing our result as in the title of this section on the retracts of $\ell_{q}$ balls for $0<q<1$ we first review the case $q\geq 1$. This is not to repeat the details available in the classical treatise of Lindenstrauss – Tzafiri [Lindstraus], we note that this is still perhaps the only book where the the main theorem about this case namely, the one stated as theorem 1.6 is dealt with. To indicate our purpose here in reviewing this case, notice that the formulation of the determining criteria for the linear retracts of $\ell^{p}$-balls as stated in the just-mentioned theorem is highly coordinate dependent, leaving a a more geometric understanding to be desired. While geometric reformulations of the same have been recorded for instance in the survey [Beata-survey], unfortunately this article neither outlines a proof nor a reference for where it could be found; the book of Lindenstrauss – Tzafiri referred therein does not give one side of the result that we need, which to the best of our knowledge is not recorded elsewhere as well. We therefore proceed to unraveling in some detail, the geometric meaning of theorem 1.6 needed for a better understanding of the linear retracts of $l^{p}$-balls.
To begin with, we observe that a linear slice of an $\ell^{p}$-ball need not be isometric to any lower dimensional $\ell^{p}$-ball. In fact, this happens only in very special circumstances, indeed precisely when the slice is a retract. As already mentioned, this result is not really new, being contained in essence in theorem 6.3 of [Beata-survey]; our purpose here is to provide the missing details, indeed a self-contained proof of this result, not to be found elsewhere. To attain this geometric result, we now recall a generalization of the notion of orthogonality in Euclidean spaces to normed spaces wherein the norm fails to be induced by any inner product; one that is most relevant for our aforementioned purposes first.
Definition 5.1.
Let $(X,\|\cdot\|)$ be a complex Banach space. For $x,y\in X$, we say that $x$ is $p$-orthogonal to $y$ and denoted $x\perp_{p}y$ if
$\begin{cases}\|x+\lambda y\|^{p}=\|x\|^{p}+|\lambda|^{p}\|y\|^{p},~\forall\lambda\in\mathbb{C},~\text{when}~p<+\infty;\\ \|x+\lambda y\|=\max\{\|x\|,|\lambda|\|y\|\},~\forall\lambda\in\mathbb{C},~\text{when}~p=+\infty.\end{cases}$
There are many notions of orthogonality and we shall not spend space in comparing the above with others except perhaps for the most popular known as Birkhoff-James orthogonality, which can be found in [Giles] and [Faulk]. Denoted suggestively by ‘BJ’, this is defined as
$x\perp_{BJ}y\Longleftrightarrow\|x+\lambda y\|\geq\|x\|~~\text{for every}~\lambda\in\mathbb{C}.$
Note that if $x\perp_{p}y$ then $x\perp_{BJ}y$ and $y\perp_{BJ}x$. Therefore, $p$-orthogonality is a stronger condition than Birkhoff-James orthogonality. This generalized orthogonality relation is, in general not symmetric. If it is indeed the case that $x\perp_{BJ}y$ and $y\perp_{BJ}x$, then we say that $x$ and $y$ are bi-orthogonal (or symmetrically orthogonal).
Note that if $u,v\in\mathbb{C}^{N}$ are mutually disjointly supported vectors then $u$ and $v$ are bi-orthogonal vectors in $(\mathbb{C}^{N},\|\cdot\|_{p})$ for $1\leq p\leq\infty$. But for each $p\neq 2$, there exists a linear subspace spanned by bi-orthogonal vectors which is not be spanned by mutually disjoint supported vectors, as illustrated in the following.
Example: Consider the linear subspace $L=f^{-1}(0)$, where the linear functional $f:\mathbb{C}^{3}\to\mathbb{C}$ defined by $f(z_{1},z_{2},z_{3})=2z_{1}+z_{2}-z_{3}$. Next we claim that $L$ is spanned by the bi-orthogonal vectors (1,1,1)(1,-1,1) and $(0,1,1)$. To prove this claim, let $v_{1}=(1,-1,1)$ and $v_{2}=(0,1,1)$. Note that
$\|v_{1}+\lambda v_{2}\|^{p}=\|(1,-1+\lambda,1+\lambda)\|^{p}=1+|1-\lambda|^{p}+|1+\lambda|^{p}$
Since $|a+b|^{p}\leq 2^{p-1}(|a|^{p}+|b|^{p})$, we obtain that $\|v_{1}+\lambda v_{2}\|^{p}\geq 1+\frac{2^{p}}{2^{p-1}}=3=\|v_{1}\|^{p}$. This shows that $v_{1}$ is orthogonal to $v_{2}$. Using similar arguments, we obtain that $v_{2}$ is orthogonal to $v_{1}$. This finishes the proof of the claim. If $L$ is spanned by mutually disjoint supported vectors then by lemma 3.2 in [Baronti-Pappini], $f$ has atmost two non-null components. Hence, we obtain that $L$ cannot be realized as a subspace spanned by mutually disjoint supported vectors.
Lemma 5.2.
Let $P$ be a linear projection map on $\mathbb{C}^{N}$. Then $\|P\|=1$ if and only if ${\rm image}(P)$ is BJ-orthogonal to $\ker(P)$.
Proof.
Assume that $P$ is a projection of norm one. We want to show that ${\rm image}(P)$ is BJ-orthogonal to $\text{ker}(P)$. Towards this, let $z\in\text{image}(P)$ and $w\in\ker(P)$; so, $P(z+\lambda w)=P(z)$ by linearity of $P$. Being a projection of norm-one $P$ is contractive, thereby for all $\lambda\in\mathbb{C}$, we have $\|z\|=\|P(z+\lambda w)\|\leq\|z+\lambda w\|$. By the notion of BJ-orthogonality recalled above, $z$ is BJ-orthogonal to $w$.
Conversely assume that $\text{image}(P)$ is BJ-orthogonal to $\ker(P)$. Since $P$ is a linear projection map, we have $\mathbb{C}^{N}=\text{image}(P)\oplus\ker(P)$. So, we can write each $z\in\mathbb{C}^{N}$ as $z=w_{1}+w_{2}$, where $w_{1}\in\text{image}(P)$ and $w_{2}\in\ker(P)$. Note that $\|P(z)\|=\|P(w_{1})\|=\|w_{1}\|$. As $w_{1}\perp_{BJ}w_{2}$, we get: $\|w_{1}\|\leq\|w_{1}+w_{2}\|=\|z\|$ i.e., $\|P(z)\|\leq\|z\|$, finishing the proof. ∎
Next, we recall the following result due to Lamperti (Corollary 2.1 of [Lamp]), which is an analogue of the parallelogram law for $\ell^{p}$-norms, determining precisely for which pair of vectors such a law holds.
Lemma 5.3.
Let $1\leq p<\infty$ with $p\neq 2$. Two vectors $u,v\in\mathbb{C}^{N}$ are mutually disjoint supported if and only if $u$ and $v$ are $p$-orthogonal.
Though the following result is not new. For instance, refer to theorem 3 of chapter 6 in [Lacey]. We therefore give a self-contained and direct exposition. For clarity, let us introduce a notation $\|\cdot\|_{p,k}$ to denote the $\ell_{p}$ norm on $\mathbb{C}^{k}$.
Proposition 5.4.
Let $W$ be a linear subspace of $(\mathbb{C}^{N},\|\cdot\|_{p,N})$. If $W$ is isometric to $(\mathbb{C}^{k},\|\cdot\|_{p,k})$ for $1\leq p\leq\infty$ with $p\neq 2$, then there exists a norm-one projection from $\mathbb{C}^{N}$ onto $W$.
Proof.
First, we consider the case $p\neq\infty$. Assume that $W$ is isometric to $(\mathbb{C}^{k},\|\cdot\|_{p,k})$. Then there exist a basis $\{u_{1},\ldots,u_{k}\}$ of $W$ and a linear isometry $S$ such that $\|u_{j}\|_{p,N}=1$ and $S(u_{j})=e_{j}$, where $e_{j}$ denotes the standard basis of $\mathbb{C}^{k}$ for $j=1,2,\ldots,k$. Note that $T:=S^{-1}$ is an isometry from $\mathbb{C}^{k}$ onto $W$ and $T(e_{j})=u_{j}$ for $j=1,2,\ldots,k$. First, we prove the result for $1\leq p<\infty$ with $p\neq 2$. Since $e_{j}$’s are disjointly supported, we can apply the above lemma 5.3 to $e_{j}$. Hence for each $j\neq l$, we have
$\|e_{j}+e_{l}\|_{p,k}^{p}+\|e_{j}-e_{l}\|_{p,k}^{p}=2(\|e_{j}\|_{p,k}^{p}+\|e_{l}\|_{p,k}^{p})$
As $T$ is an isometry, we obtain that
$\|T(e_{j})+T(e_{l})\|_{p,N}^{p}+\|T(e_{j})-T(e_{l})\|_{p,N}^{p}=2(\|T(e_{j})\|_{p,N}^{p}+\|T(e_{l})\|_{p,N}^{p})$
Again applying the above lemma 5.3, we get that the $u_{j}$’s are mutually disjoint supported vectors in $\mathbb{C}^{N}$. Following the proof of proposition 2.a.1 in [Lindstraus], the details regarding the construction of norm-one projection are given as follows. For each $j=1,2,\ldots,k$, let $\sigma_{j}:=supp(u_{j})$. Consider the linear functional $f_{j}:W\to\mathbb{C}$ defined by $f_{j}(u_{l})=\delta_{jl}$ (where $\delta_{jl}$ denotes the Kronecker-delta function) and extended to other vectors of $W$ by linearity.
First we want to prove that $\|f_{j}\|_{p,N}=1$ for each $j=1,2,\ldots,k$. To prove this, write each $z\in W$ as $z=\sum_{l=1}^{k}c_{l}u_{l}$ for some scalars $(c_{l})_{l=1}^{k}$. Note that
(5.1) $|f_{j}(z)|=\left|f_{j}\left(\sum_{l=1}^{k}c_{l}u_{l}\right)\right|=\left|\sum_{l=1}^{k}c_{l}f_{j}(u_{l})\right|=|c_{j}|\leq\|(c_{1},\ldots,c_{k})\|_{p,k}=\|S(z)\|_{p,k}.$
Since $S$ is an isometry, we obtain that $|f_{j}(z)|\leq\|z\|_{p,N}$. This implies that $\|f_{j}\|_{p,N}\leq 1$ for all $j=1,2,\ldots,k$. Since $f_{j}(u_{j})=1$, we obtain that $\|f_{j}\|_{p,N}=1$. By the Hahn – Banach extension theorem, $f_{j}$ can be extended to a linear functional on $\mathbb{C}^{N}$ of the same norm which we also denote by $f_{j}$. So, $f_{j}:\mathbb{C}^{N}\to\mathbb{C}$ is a linear functional with $\|f_{j}\|_{p,N}=1$. For each $j$ with $1\leq j\leq k$, write $f_{j}(z)=\sum_{i=1}^{N}c_{j,i}z_{i}$ for some scalars $(c_{j,i})_{i=1}^{N}$. Define another linear functional $u_{j}^{*}$ on $\mathbb{C}^{N}$ by $u_{j}^{*}(z)=\sum_{i\in\sigma_{j}}c_{j,i}z_{i}$. Note that
$\|u_{j}^{*}\|_{p,N}^{q}=\sum_{i\in\sigma_{j}}|c_{j,i}|^{q}\leq\sum_{i=1}^{N}|c_{j,i}|^{q}=\|f_{j}\|_{p,N}^{q}=1$
where $q$ is the dual exponent of $p$ (i.e. $1/p+1/q=1$.). This implies in particular that $|u_{j}^{*}(\tilde{e_{i}})|=|c_{j,i}|\leq 1$, where $\tilde{e_{i}}$ denote the standard basis of $\mathbb{C}^{N}$. As $u_{j}^{*}(u_{j})=1$, we obtain that $\|u_{j}^{*}\|_{p,N}=1$ for every $j$. Now we define the projection $P:\mathbb{C}^{N}\to W$ by $P(z)=\sum_{j=1}^{k}u_{j}^{*}(z)u_{j}$. To prove this proposition, it suffices to show that $\|P\|_{p,N}=1$. If $z=\sum_{i=1}^{N}a_{i}\tilde{e_{i}}$ then $u_{j}^{*}(z)=\sum_{i=1}^{N}a_{i}u_{j}^{*}(\tilde{e_{i}})=\sum_{i\in\sigma_{j}}a_{i}c_{j,i}$. Since $\|u_{j}^{*}\|_{p,N}=1$, $|u_{j}^{*}(z)|^{p}\leq\sum_{i\in\sigma_{j}}|a_{i}|^{p}$. Hence we obtain that $\|P(z)\|_{p,N}^{p}=\sum_{j=1}^{k}|u_{j}^{*}(z)|^{p}\leq\|z\|^{p}$. This proves that $P$ is a norm-one projection from $\mathbb{C}^{N}$ onto $W$.
Next, we prove the result for $p=\infty$. Write $S(z)=(S_{1}(z),\ldots,S_{k}(z))$. Note that $S_{j}(z)$ is a linear functional on $W$ for each $j$ with $1\leq j\leq k$. By the Hahn – Banach theorem, $S_{j}$ can be extended to a linear functional on $\mathbb{C}^{N}$ of the same norm which we denote by $\tilde{S_{j}}$. So, $\tilde{S_{j}}:\mathbb{C}^{N}\to\mathbb{C}$ is a linear functional so that $\|\tilde{S_{j}}\|=\|S_{j}\|\leq 1$. Define the linear map $\tilde{S}:\mathbb{C}^{N}\to\mathbb{C}^{k}$ by $\tilde{S}(z)=(\tilde{S}_{1}(z),\ldots,\tilde{S}_{k}(z))$. Note that $\|\tilde{S}\|\leq 1$ because $\|\tilde{S}\|=\sup\limits_{\begin{subarray}{c}1\leq j\leq N,~\|z\|=1\end{subarray}}|S_{j}(z)|\leq 1$. Since $\tilde{S}(u_{j})=e_{j}$, we get that $\|\tilde{S}\|=1$. Consider the linear map $P:\mathbb{C}^{N}\to W$ defined by $P(z)=T\circ\tilde{S}(z)$. Note that $P$ maps $\mathbb{C}^{N}$ onto $W$ and $P_{|_{W}}={\rm id}_{W}$. Hence, we get that $P$ is a norm one projection from $\mathbb{C}^{N}$ onto $W$. ∎
Proposition 5.5.
Let $\mathbb{B}_{p}$ be the $\ell_{p}$-ball in $\mathbb{C}^{N}$ where $1\leq p\leq\infty$. The linear subspaces of $\mathbb{C}^{N}$ such that $(\mathbb{B}_{p})_{L}:=\mathbb{B}_{p}\cap L$ is a retract of $\mathbb{B}_{p}$ are precisely those for which $(\mathbb{B}_{p})_{L}$ is biholomorphic to a lower dimensional $\ell_{p}$-ball $\mathbb{B}_{p}^{k}$ for some $k$ with $1\leq k\leq N$.
Proof.
Assume that $(\mathbb{B}_{p})_{L}$ is biholomorphic to a lower dimensional $\ell_{p}$-ball $\mathbb{B}_{p}^{k}\subset\mathbb{C}^{k}$. By the theorem of W. Kaup and H.Upmeier in [Kaup-Upm] (exposited as Remark 16.4.2 d) in [Jrncki_invrnt_dst]), after passing to a different biholomorphic mapping we may assume that $\varphi(0)=0$. By an application of Cartan’s theorem, we get that $\varphi$ is linear. Note that for each $v\in(\mathbb{B}_{p})_{L}$, we have
$\|v\|_{p}=h_{(\mathbb{B}_{p})_{L}}(v)=K_{(\mathbb{B}_{p})_{L}}(0,v)=K_{\mathbb{B}_{p}^{k}}(0,\varphi(v))=\|\varphi(v)\|_{p,k}$
This implies that $\varphi$ is a linear isometry from $L$ onto $(\mathbb{C}^{k},\|\cdot\|_{p,k})$. By the above proposition 5.4, we obtain that $(\mathbb{B}_{p})_{L}$ is a linear retract of $\mathbb{B}_{p}$.
Conversely assume that $(\mathbb{B}_{p})_{L}$ is a linear retract of $\mathbb{B}_{p}^{N}$. First we consider the case $1\leq p<\infty$. By theorem 2.a.4 in [Lindstraus], there exists an isometry from $L$ to $(\mathbb{C}^{k},\|\cdot\|_{p,k})$. This implies in particular that $(\mathbb{B}_{p})_{L}$ is biholomorphic to $\mathbb{B}_{p}^{k}$ for some $k$ with $1\leq k\leq N$. Next we consider the case for $p=\infty$. By proposition 3.16, there exists an isometry from $L$ to $(\mathbb{C}^{k},\|\cdot\|_{\infty})$. This proves that $\Delta^{N}_{L}$ is biholomorphic to lower dimensional polydisk in $\mathbb{C}^{k}$. ∎
We conclude the characterizations of retracts of $l^{p}$-balls in the convex case $p\geq 1$.
@@
A linear subspace $R$ of the $\ell_{p}$-ball $\mathbb{B}_{p}$ is a $k$-dimensional retract of the $\mathbb{B}_{p}$ in $\mathbb{C}^{N}$ if and only if there exist a finite sequence of retracts $R_{1}\subset R_{2}\subset\ldots\subset R_{k}\subset\mathbb{B}_{p}$ such that for each $j=1,2,\ldots,k$, $R_{j}$ is a $j$-dimensional linear retract of $\mathbb{B}_{p}$ and $R_{k}=R$.
@@
As the sufficiency part is utterly trivial, we need assume that $R$ is a linear retract of the $\ell_{p}$-ball $\mathbb{B}_{p}$ and verify the reachability of $R$ via a monotonically increasing (finite) sequence of retracts as asserted. By theorem 2.a.4 in [Lindstraus], $R$ is isometric to $k$-dimensional $\ell_{p}$-ball $\mathbb{B}_{p}^{k}$. Recalling there always exist retracts of all possible dimensions in $\mathbb{B}_{p}$, we have in-particular, a (k1)(k-1)-dimensional retract $R_{k-1}$ of $\mathbb{B}_{p}^{k}$. By lemma 2.13, $R_{k-1}$ is a linear retract of $\mathbb{B}_{p}$. Again applying theorem 2.a.4 in [Lindstraus] and lemma 2.13 to $\mathbb{B}_{p}^{k-1}$, we obtain that there exists a (k2)(k-2)-dimensional retract $R_{k-2}$ of $\mathbb{B}_{p}$. Continuing this process, until we get a one-dimensional retract $R_{1}$ of $\mathbb{B}_{p}$, renders the required sequence as in the statement of the theorem. ∎
@@
We now put the above result together with the foregoing more geometric formulation of the linear retracts of $\ell_{p}$-balls as being precisely those linear intercepts which are isometric copies of lower dimensional $\ell_{p}$-balls. Doing so, leads to the conclusion that for any given retract $R$, there is a decreasing chain of isometric copies of lower dimensional $\ell^{p}$-balls reaching $R$ as in the above theorem. This is a feature shared by the $\ell_{p}$-‘balls’ even for positive values of the parameter $p<1$. Indeed, we now move onto this case dealt with by theorem 1.8. Before going into its proof, we need to make a couple of observations about the geometry of the $\ell_{q}$-ball $D_{q}$ in $\mathbb{C}^{N}$ for $q<1$ – for a vivid distinction between the convex case $p\geq 1$ with the non-convex case $p<1$, we use the symbol $q$ rather than $p$, for the latter case. To begin with then, although the absolute image of $D_{q}$ is concave (easily and vividly seen in the two-dimensional case), $D_{q}$ is actually not concave (nor convex) domain in $\mathbb{C}^{N}$; not even locally, near any of the smooth boundary points. We may establish this by computing the real Hessian at any smooth boundary point with positive real coordinates as $D_{q}$ is multicircular; we show this in the two-dimensional setting for simplicity, while it holds in higher dimensions as well by straight-forward extension of the arguments and computations. Rewrite the defining function for $D_{q}$ as
@@
$H$ $:=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial z\partial\overline{w}}\\ \frac{\partial^{2}\varphi}{\partial w\partial\overline{z}}&\frac{\partial^{2}\varphi}{\partial w\partial\overline{w}}\end{bmatrix}=\begin{bmatrix}\frac{q^{2}}{4}|z|^{q-2}&0\\ 0&\frac{q^{2}}{4}|w|^{q-2},\end{bmatrix}$
$\tilde{H}$ $:=$ $\begin{bmatrix}\frac{\partial^{2}\varphi}{\partial z^{2}}&\frac{\partial^{2}\varphi}{\partial z\partial w}\\ \frac{\partial^{2}\varphi}{\partial w\partial z}&\frac{\partial^{2}\varphi}{\partial w^{2}}\end{bmatrix}=\begin{bmatrix}\frac{q}{2}(\frac{q}{2}-1)|z|^{q}z^{-2}&0\\ 0&\frac{q}{2}(\frac{q}{2}-1)|w|^{q}w^{-2}\end{bmatrix}$
@@
(5.2) $vH\overline{v}^{T}_{|_{(x,y)}}+Re(vHv^{T})_{(x,y)}\\ =\frac{q^{2}}{4}\left(x^{q-2}|v_{1}|^{2}+y^{q-2}|v_{2}|^{2}\right)+\frac{q}{2}(\frac{q}{2}-1)Re\left(x^{q-2}v_{1}^{2}+y^{q-2}v_{2}^{2}\right).$